There is a cross section of a clear spherical container with some fruit juice in it. The radius of the sphere is 18cm and the perpendicular distance from the centre of the sphere to the surface of the fruit juice is 9cm, calculate the diameter of the surface of the fruit juice in the sphere.
Answer by RAY100(1637)

The sphere, liquid, and radius, makes a right triangle, with a hypotenuse equal to the spherical radius of 18 cm, and one leg of length 9 cm.
.
Using the pythagorean theorem, the radius of the liquid can be found.
.
c^2 = a^2 +b^2,,,,or 18^2 = 9^2 + r^2,,,,r^2 = 18^2 -9^2 = 243,,,,r=15.588 cm
.
or the diameter = 2*15.588 = 31.177 cm
.